OpenStudy (anonymous):

tan theta+cot theta= 1/sin theta cos theta

OpenStudy (raden):

remember the identities that : tanθ = sin θ /cos θ and cot θ = cos θ /sin θ

OpenStudy (jdoe0001):

what would be the identities equivalents for TAN and COT in terms of cosine and sine?

OpenStudy (anonymous):

okay...

OpenStudy (jdoe0001):

as RadEn just pasted, those are it, just ADD them up, get the LCD and add them up, see what you get

OpenStudy (anonymous):

so the answer is what RadEn said?

OpenStudy (jdoe0001):

well, no, replace tangent and cotangent, on the left-side, by what he said :)

OpenStudy (jdoe0001):

$$ tan(\theta) = \cfrac{sin(\theta)}{cos(\theta)}\ \ \&\ \ \ cot(\theta) = \cfrac{cos(\theta)}{sin(\theta)}\ thus\\ \cfrac{sin(\theta)}{cos(\theta)}+\cfrac{cos(\theta)}{sin(\theta)} = \cfrac{}{\boxed{LCD?}} $$

OpenStudy (raden):

well, we have to prove the LHS can becomes like RHS tanθ +cot θ = sin θ /cos θ + cos θ /sin θ = (sin^2 θ + cos^2 θ)/sinθcosθ then use the identity : sin^2 θ + cos^2 θ = 1 so, = 1/sinθcosθ